What is the solution for the exponential equation 9^-x=(1/3)
Answer by lwsshak3(11628)

What is the solution for the exponential equation 9^-x=(1/3)
9^-x=1/3
(3^2)^-x=3^-1
3^-2x=3^-1
-2x=-1
x=1/2